OpenStudy (anonymous):

check my answer? simplify the expression; (-5i)(-3i) this would be 15i since a - times a - is a positive right?

OpenStudy (math&ing001):

A - times a - is indeed a positive, but i*i is not equal to i

OpenStudy (anonymous):

oh wait, I is equal to -1. so that would make it -15i wouldn't it?

OpenStudy (math&ing001):

i*i=i^2=-1

OpenStudy (anonymous):

That'll be 15i to the second power, since you would add the exponents.

OpenStudy (anonymous):

15i^2 isn't an answer choice... it would be -15i wouldn't it?

OpenStudy (math&ing001):

It would be -15

OpenStudy (anonymous):

Wait, what type of math do you have? Because that's the right answer; you have to add the exponents and multiply like bases.

OpenStudy (anonymous):

I'm in Algreba 101 and that seems to be the correct one...

OpenStudy (anonymous):

its a compound inequality so its different, and im not sure, because 15i^2 wasn't one of the answers. but yeah, I would think that your right. that's not one of the answers though so idk

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

thanks for your input though!

OpenStudy (anonymous):

You're welcome!

OpenStudy (math&ing001):

@Walmart789456 those are called imaginary numbers which are defined by the property i^2=-1